How much energy is required to move hydrogen electron from 1st to 2nd shell


1
Expert's answer
2021-03-29T06:08:46-0400

1/λ = R( 1/ N1^2 - 1/N2^2)

Where n< n2

In this case n2=∞

 so this becomes:

1/λ=R/1^2

λ = 1/R

correct the units of the Rydberg Constant to m^−1

R=1.097 × 10^ -7 m^-1

Therefore the wavelength of the photon required to eject the electron will be given by

λ= 1/1.097×10 ^-7 = 9..1157×10^−8m

E = hf = hc/λ =

=6.626×10^−34×3×10^8/9.1157×10^−8

= 2.180 × 10^-18 J

E=2.180×10^−18×6.02×10^23=

= 13.1×10^5 J/mol/ 1000

E=1310kJ/mol